OpenStudy (anonymous):

a fully loaded truck is carrying grain to a storage silo. If the truck is 8 ft wide 40 ft long and 4 ft deep how many bushels of grains is it carrying? (1 bushel = 1.25 cubic feet)

OpenStudy (anonymous):

Volume of grain the truck is carrying = 8 * 40 * 4 = 1280 cubic feet No of bushels it is carrying = 1280 / 1.25 = 1024 bushels.